OpenStudy (anonymous):

A motorcycle can travel 60 miles per gallon. Approximately how many gallons of fuel will the motorcycle need to travel 40 km? [1 mile = 1.6 km] 0.04 0.08 0.20 0.42

OpenStudy (danjs):



OpenStudy (danjs):

40 km at 1 mile is 1.6 km and 60 miles to 1 gallon

OpenStudy (anonymous):

oh

OpenStudy (danjs):

\[\frac{ 40 ~km }{ 1 }*\frac{ 1~mi }{ 1.6~km }*\frac{ 1~gallon }{ 60~mi } = \]

OpenStudy (danjs):

notice the units km and mi, cancel and you are left with gallons

OpenStudy (anonymous):

sooo how many gallons would tht be?

OpenStudy (danjs):

40 / (1.6*60) = 40 / 96 = about 0.41666 or 0.42

OpenStudy (anonymous):

ok thank youu!!